A worker requires 8 tanks of crude oil and petrol to get a work done. He gains 5000 

Ghana cedis per tank of crude oil used and 3000 Ghana cedis per petrol used for any 

work done. To finish the work, he further requires 10 tanks by combining 2 tanks of 

crude oil and a tank of petrol. What is the maximum profit he can make?


1
Expert's answer
2021-03-30T07:16:13-0400

"\text{Let x = no of tanks of crude oil and y = no of tanks of petrol. }"

"\text{We are to maximise }5000x +3000y\text{ subject to conditions:}"

"2x+y=10-(1)"

"x+y =8-(2)"

"\text{Using substitution method}"

x=2 and y=6

"\text{Substituting x and y in 5000x + 3000y, we have that the maximum profit he can make }"

"\text{is 28000 Ghana cedis}"